### v2

# Analytic bulk-edge connection in circular-symmetric models

Abstract:
We propose a systematic analysis of the eigenfunctions of two-band systems in two dimensions with a circular edge. Our approach is based on an analytic continuation of the wavenumber, which yields a mapping from the bulk modes to the edge modes. Phase relations of the eigenfunctions are described by their mapping onto a three-dimensional field of unit vectors. This mapping is studied in detail for a two-band Laplacian model and a Dirac model. The direction of the unit vector identifies the phase relation of the eigenfunctions and enables us to distinguish between the upper band, the lower band and the edge spectrum. Bulk and edge modes are spectrally separated, which results in two transitions from delocalized bulk modes to localized edge modes. These transitions are accompanied by transitions of the phase relations. Our analytic approach is compared with the topological bulk-edge correspondence, which is based on the Chern number of the bulk.
1 Introduction
Recent studies on complex photonic and phononic systems have opened up new perspectives for analyzing wavefunctions, which are solutions of the Schrödinger or Dirac equations of single quantum particles [1, 2, 3, 4]. Besides the spectrum, wavefunctions carry crucial information about both classical and quantum systems, particularly regarding their topological properties, which influence optical and transport behaviors. In contrast to electronic systems, wavefunctions can be directly observed in classical systems, and their coherence is often easier to manipulate than in electronic systems. For example, the intensity of an electromagnetic or sound wave, represented by the magnitude of the wavefunction, can be measured locally at position $\bf r$}. Moreover, the structure of the two-component wavefunction of a two-band Hamiltonian offers even deeper insights, since the relative phase between the two components plays a central role in determining the system’s properties.
In a system with edges we can distinguish bulk and edge modes. Both are eigenfunctions of the Hamiltonian but for different eigenvalues and with qualitatively different properties: the bulk modes are wave-like functions, extended over the entire systems, while the edge modes decay exponentially away from the edge. Edge modes play a crucial role in the quantum Hall effect [5, 6]. They are quite robust but at the same time they are also sensitive to the boundary conditions. Several types of boundary conditions and their corresponding edge modes were studied in electromagnetic systems [7].
The goal of this work is to analyze the connection between those two types of modes, which is related to the bulk-edge connection. A similar idea to describe the bulk-edge correspondence was pursued for topological materials, based on the connection of topological invariants of the bulk (Chern numbers) and the number of edge modes. In particular, this has been discussed intensively in the context of the quantum Hall effect and the 2D Dirac equation, pioneered by the work of Hatsugai [8], and followed by a broader application of this idea in Refs. [9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22].
In the following we will develop a more direct approach for the bulk-edge connection, using an analytic continuation without relying on topological invariants. For this purpose we consider the equation $H\Psi_{E}=E\Psi_{E}$ with $\Psi_{E}=(\psi_{E,1},\psi_{E,2})$ for some models with a two-band Hamiltonian $H$, which is assumed to act on a two-dimensional space and has the general structure
$H=h_{1}\sigma_{1}+h_{2}\sigma_{2}+h_{3}\sigma_{3}\equiv{\vec{h}}\cdot{\vec{\sigma}}$ (1)
in terms of Pauli matrices $\sigma_{j}$ with ${\vec{\sigma}}=(\sigma_{1},\sigma_{2},\sigma_{3})^{T}$. We focus on a circular symmetry, which is easy to realize experimentally.
The analytic continuation of wavefunctions is a powerful method, which has been applied to many physical problems. A typical example is the scattering theory, where bound states are found as poles of the scattering matrix or the Green’s function. In this paper we will solve the eigenvalue problem $H\Psi_{E}=E\Psi_{E}$ for the bulk modes and construct the corresponding edge modes by an analytic continuation, extending an idea employed to a 2D superconductor [23].
2 Two-band models
Inspired by the electronic properties of graphene [24] as well as by wave properties in photonic and phononic systems on a honeycomb structure, the 2D Dirac Hamiltonian with ${\vec{h}}=(i\partial_{x},i\partial_{y},m)$ with Dirac mass $m$ in the Hamiltonian of Eq. (1) has been extensively studied in recent years [1, 2, 3, 4, 25]. The eigenfunction of a translational-invariant two-band Hamiltonian reads
[…]
where ${\bf r}=(x,y)$ is the position in space and $\bf k$} is the wavevector. Moreover, in a circular symmetric system the position is parametrized by the radius $r$ and the polar angle $\alpha$, and the two components of the wavevector are parametrized by the wavenumber $k$ and the angular momentum number $n=0,\pm 1,\ldots$ such that the eigenfunction reads
[…]
2.1 Topological properties
Topological properties of the two-component wavefunction $\Psi({\bf r})=(\psi_{1},\psi_{2})^{T}$, such as its chirality, can be analyzed through topological invariants, such as winding numbers or Chern numbers. Another interesting quantity in this context is the Hermitian tensor field $\psi_{i}^{*}({\bf r})\psi_{j}({\bf r})$, which is gauge-invariant in the sense that phase factors of the wavefunctions cancel each other in the product and only phase differences survive. Using $\psi_{j}=|\psi_{j}|e^{i\varphi_{j}}$, we define a real three-dimensional vector $\vec{s}$} from the Hermitian tensor $\psi_{i}^{*}({\bf r})\psi_{j}({\bf r})$ as
$s_{1}=\frac{2|\psi_{1}||\psi_{2}|}{|\psi_{1}|^{2}+|\psi_{2}|^{2}}\cos(\varphi_{2}-\varphi_{1})\ ,\ \ s_{2}=\frac{2|\psi_{1}||\psi_{2}|}{|\psi_{1}|^{2}+|\psi_{2}|^{2}}\sin(\varphi_{2}-\varphi_{1})\ ,\ \ s_{3}=\frac{|\psi_{1}|^{2}-|\psi_{2}|^{2}}{|\psi_{1}|^{2}+|\psi_{2}|^{2}},$ (4)
which characterizes the eigenfunctions according to the magnitudes of their vector components and their phase differences. The vector components $s_{1},s_{2}$ provide a winding number of the wavefunctions through their phase dependence. Therefore, this gauge-invariant field is reminiscent of the Berry connection. Direct inspection reveals that ${\vec{s}}=(s_{1},s_{2},s_{3})^{T}$ is a three-dimensional unit vector due to $s_{1}^{2}+s_{2}^{2}+s_{3}^{2}=1$. Thus, the trajectory of ${\vec{s}}({\bf r})$ is a horizontal circle on the unit sphere when we vary the phase difference $\varphi_{2}-\varphi_{1}$ from 0 to $2\pi$, as visualized in Fig. 1. The mapping $(\psi_{1}({\bf r}),\psi_{2}({\bf r}))^{T}\to{\vec{s}}({\bf r})$, which reflects a mapping of the two-dimensional plane to the unit sphere $S^{2}$, enables us to identify the wavefunction as a structure on a compact manifold. An expansion of $\vec{s}$} in terms of Pauli matrices yields
${\vec{s}}=\frac{\Psi\cdot{\vec{\sigma}}\Psi}{\Psi\cdot\Psi}.$ (5)
In general, the mapping ${\vec{h}}\to{\vec{s}}$ is central for a two-band Hamiltonian. It will be shown subsequently that ${\vec{s}}({\bf r})$ characterizes the properties of the eigenfunctions through its trajectories when we vary the position $\bf r$}. Besides $\vec{s}$}, the local intensity or signal strength $I=\Psi\cdot\Psi=|\psi_{1}|^{2}+|\psi_{2}|^{2}$ is another relevant quantity to characterize the eigenfunctions of $H$, which yields the spatial distribution of the signal strength. The distribution of bulk modes is quite different from the distribution of edge modes, since for the latter it is concentrated only at the edge(s). In classical systems the spatial integral of $I$ is the energy stored in the sample, while for quantum systems it is 1.
Figure 1: Sphere of the $S^{2}$ field $\vec{s}$}.
Topological properties, such as the chirality, can be identified, for instance, with the edge vorticity (EV) of a circular edge, which is defined as
${\cal V}=\int_{\cal F}[\nabla\times{\vec{s}}]\cdot d^{2}{\bf r},$ (6)
[…]
The phase difference in Eq. (4) depends on the location $(r,\alpha)$, and $\nabla\times{\vec{s}}$ can identify a vortex. A simple example is the eigenfunction of the translational-invariant 2D Dirac Hamiltonian:
[…]
for the eigenvalues $\pm\sqrt{m^{2}+k^{2}}$. This gives immediately a spatial uniform ${\vec{s}}_{\bf k}=2\rho_{k;\pm}(\cos\gamma,\sin\gamma,0)^{T}$ with $\gamma={\rm arg}(k_{x}+ik_{y})$, while the divergence with respect to $\bf k$} is proportional to $\rho_{k;\pm}$. This indicates a source (sink) of ${\vec{s}}_{\bf k}$ at the center for the lower (upper) band for $m>0$. On the other hand, the EV vanishes due to ${\cal V}=0$. $\rho_{k;\pm}$ switches its sign with $m\to-m$. In the following we will assume that $m>0$.
3 Topological bulk-edge correspondence and analytic bulk-edge connection
Almost all experimental samples have at least one edge, where a typical example is a disk with a circular edge. The contribution of the edge to the properties of the sample is exponentially small and is usually ignored in the macroscopic description. This changed after the discovery of the quantum Hall effect with a robust Hall conductivity [26], where the latter was attributed to edge modes [5]. It soon turned out that it is crucial to understand how the edge modes are related to the bulk modes, since the existence of edge modes inside the spectral gap of the bulk modes plays a crucial role to explain the transport properties of the quantum Hall effect [8]. A constructive approach to the bulk-edge correspondence can be based on the projection in $\bf k$} space perpendicular to a straight edge [9, 27]. The corresponding set of one-dimensional solutions, parametrized by the wavevector along the edge, provides the edge modes and their energies. This approach is rather involved though. In particular, it requires the solution of the one-dimensional equations for the edge. Therefore, we suggest here a different approach, based on an analytic continuation of the wavevector, which we applied previously to the Bogoliubov de Gennes equation for superconducting double layers [23]. To distinguish it from the topological bulk-edge correspondence (TBEC), it will be called subsequently analytic bulk-edge connection (ABEC). An advantage of this approach is that it is sufficient to solve only the bulk equation. We will argue in the following that edge modes appear inside a spectral gap in systems without specific reference to topology of the eigenfunctions (i.e. it is applicable also to eigenfunctions with vanishing EV). It will be shown that their origin is purely geometric, typically due to an edge. They are specific for a given Hamiltonian $H$ but can be determined within a systematic approach, which is based on an analytic continuation of the bulk modes.
Starting point is the spectrum $E(k)=(m^{2}+k^{2})^{1/2}$ of the 2D Dirac Hamiltonian with real wave number $k$ and a positive Dirac mass $m$. This example is considered for its simplicity but the concept can be directly generalized to other Hamiltonians with the bulk spectrum $E({\bf k})=(m^{2}+h({\bf k}))^{1/2}$ with an analytic function $h$. The square root provides two bands $\pm\sqrt{m^{2}+k^{2}}$ for the 2D Dirac case, which are separated by a gap $2m$. Formally, this spectrum can be extended by continuing the real wave number $k$ to the complex plane. This creates a connected Riemann surface (cf. Fig. 2a,c). Therefore, the gap is bridged by modes with complex $k$. However, a complex eigenvalue $E$ is not physical, such that we must restrict the analytic continuation to real values of $E$. This is the case for real $k$ and for purely imaginary values $k=ic$ with $m\leq c\leq m$, where the latter gives $E=\pm\sqrt{m^{2}-c^{2}}$. This means that the spectrum separates in a bulk spectrum with $E\geq m$ and $E\leq-m$ and an edge spectrum with m<E<m-m<E<m.
For the analytic continuation of the eigenfunctions we assume that the bulk mode $\Psi(x,y)$ is either a plane wave $e^{ik_{1}x+ik_{2}y}$ for a translational-invariant system or a circular wave $e^{ikr+in\alpha}$ with polar coordinates $(r,\alpha)$ for a circular-invariant system. Then we obtain from the analytic continuations $k_{j}\to ic_{j}$ and $k\to ic$ either $e^{-c_{1}x-c_{2}y}$ or $e^{-cr+in\alpha}$, respectively. These modes decay exponentially for $c_{1}x>0$, $c_{2}y>0$, representing a so-called corner mode, or $cr>0$, representing a circular edge mode. The boundary conditions play obviously a crucial role here, since the edge modes should decay exponentially from the edge. This selects whether $_{j}, cc$ are positive or negative. For a large disk with a central hole, for example, this means that $c>0$ at the edge of the hole.
We note that the analytic continuation does not affect the Hamiltonian, only the spectrum and the eigenfunctions. The analytic continuation can be reversed such that we get the bulk modes from the edge modes. This can be useful when we observe or manipulate the edge modes and want to determine the corresponding bulk modes.
4 Special examples
4.1 Single-band Laplacian model
We consider the eigenvalue problem $H\Psi_{E}=E\Psi_{E}$ with the Hamilton operator $H$, acting on a 2D space, and the real energy eigenvalue $E$. This equation appears also in many areas of classical physics, for instance, in microwave systems, photonics and phononics [1, 2, 3, 4]. In the following, our models are characterized by different Hamiltonians $H$. A prototype of $H$ is a single Laplacian $H=-\Delta$ on a disk with a non-negative bulk spectrum. For a the circular geometry $\Delta$ is parametrized by polar coordinates $(r,\alpha)$ as
$\Delta=\frac{\partial^{2}}{\partial r^{2}}+\frac{1}{r}\frac{\partial}{\partial r}+\frac{1}{r^{2}}\frac{\partial^{2}}{\partial\alpha^{2}}.$ (8)
The Bessel functions $J_{n}(kr)$ and $Y_{n}(kr)$, multiplied by a phase factor $e^{in\alpha}$, are eigenfunctions of this Laplacian with eigenvalues $k^{2}$ [28]. Two linearly independent solutions are given by the linear combinations
$\phi^{\pm}_{k,n}(r,\alpha):=J_{n}(kr)\pm iY_{n}(kr)=c_{n}\frac{e^{\pm ikr}}{\sqrt{kr}}+O\left(\frac{1}{kr}\right)\ \ (n=0,\pm 1,\ldots)$ (9)
with $c_{n}=e^{\mp i(\pi n+\pi/2)/2}\sqrt{2/\pi}$. The eigenvalues $k^{2}$ of $\Delta$ are independent of $n$ due to the rotational invariance of the Laplacian.
The analytic continuation $k\to ic$ with a real $c$ yields negative eigenvalues $k^{2}\to-c^{2}$ and the modified Bessel functions as (cf. App. B)
$J_{n}(kr)\to J_{n}(icr)=e^{i\pi n/2}I_{n}(cr)\ ,\ \ J_{n}(kr)+iY_{n}(kr)\to H_{n}^{(1)}(icr)=-\frac{2i}{\pi}e^{-i\pi n/2}K_{n}(cr),$ (10)
where the modified Bessel function $I_{n}(cr)$ ($K_{n}(cr)$) increases (decreases) exponentially for $cr>0$ and $K_{0}(cr)$ diverges for $cr\to 0$. A proper linear combination gives a unique solution that satisfies the boundary conditions at the edge. For a disk this is matched by $I_{n}(cr)$ with a finite value at $r=0$ and for a hole on a large disk it is matched by $K_{n}(cr)$. The exponential decay rate of the edge modes is given by $1/|c|=1/\sqrt{|E|}$, indicating a radial shrinking of the edge modes with decreasing energy $E=-c^{2}$.
The requirement of the analytic continuation is that (i) the eigenvalues remain real and (ii) edge modes decay exponentially from an edge (i.e., they are evanescent modes). Thus, they depend strongly on the boundary conditions. We conclude that the spectrum of $\Delta$ consists of two branches, which are parametrized by the complex wavenumber $k$. One branch is for the bulk modes with $E\geq 0$ and the other is for the edge modes with $E<0$. Although both spectra are real, the bulk spectrum is parametrized by a real wavenumber, while the edge spectrum is parametrized by an imaginary wavenumber.
4.2 Two-band Laplacian model
The 2D Laplacian $\Delta$ can be used to construct the Hamiltonian in Eq. (1) with $h_{1}=-\Delta$, $h_{2}=0$ and $h_{3}=m$. This yields the real symmetric matrix
[…]
Since the three-component vector ${\vec{h}}=(-\Delta,0,m)$ is only a one-dimensional line in 2D, the bands have vanishing Chern numbers.
Now we consider $H_{\rm 2\Delta}$ on a space with a circular geometry (e.g., a disk). Using the Laplacian in Eq. (8) and its eigenfunctions in Eq. (9), the ansatz
[…]
with the wavenumber $k$ yields
[…]
with the eigenvalues $E_{k}=\pm\sqrt{m^{2}+k^{4}}$. The components of the eigenvectors $(a_{1},a_{2})^{T}$ read
$a_{1}=k^{2}\ ,\ \ a_{2}=m\pm\sqrt{m^{2}+k^{4}}$ (14)
with an additional normalization that depends on the specific physical system. The $k^{4}$ dependence of the eigenvalues suggests the analytic continuation $k\to c=e^{\pm i\pi/4}c$ with $\sqrt{m}\leq c\leq\sqrt{m}$, such that $E_{k}=\pm\sqrt{m^{2}+k^{4}}\to\pm\sqrt{m^{2}-c^{4}}$ is real. In Fig. 2c the analytic continuation of the bulk spectrum $E_{k}=\pm\sqrt{m^{2}+k^{4}}$ is depicted as a Riemann surface and compared with the Riemann surface of $E_{k}=\pm\sqrt{m^{2}+k^{2}}$ in Fig. 2a. It should be noted that this analytic continuation creates a non-Hermitian matrix in Eq. (13):
[…]
There is no conflict within our approach though, since we calculate the eigenfunctions before the analytic continuation, where the matrix is real symmetric. Then the analytic continuation of the eigenfunctions and their inner product yields the correct result for the non-Hermitian matrix, as it can be directly checked for this simple case.
Figure 2: The real part of $E(k)$ for a complex wavenumber $k$. a) Riemann surface of the real part of $E(k)=(m^{2}+k^{2})^{1/2}$ and b) the related bulk and edge spectra for a real $E(k)$. c) Riemann surface of the real part of $E(k)=(m^{2}+k^{4})^{1/2}$.
The analytic continuation $k\to e^{\pm i\pi/4}c$ of the eigenfunctions yields
$\phi^{+}_{n}(e^{\pm i\pi/4}cr)=J_{n}(e^{\pm i\pi/4}cr)+iY_{n}(e^{\pm i\pi/4}cr)=c_{n}e^{\mp cr/\sqrt{2}}e^{icr/\sqrt{2}}\left[\frac{e^{\mp i\pi/8}}{\sqrt{cr}}+O\left(\frac{1}{cr}\right)\right]$
$\phi^{-}_{n}(e^{\pm i\pi/4}cr)=J_{n}(e^{\pm i\pi/4}cr)-iY_{n}(e^{\pm i\pi/4}cr)=c_{n}e^{\pm cr/\sqrt{2}}e^{icr/\sqrt{2}}\left[\frac{e^{\mp i\pi/8}}{\sqrt{cr}}+O\left(\frac{1}{cr}\right)\right].$ (16)
These functions grow (decay) exponentially with $cr>0$, where boundary conditions select the unique solution. For instance, $\phi^{+}(e^{-i\pi/4}cr)$ and $\phi^{-}(e^{i\pi/4}cr)$ are valid as an edge mode inside a disk with radius $r_{0}<\infty$, where the mode decays exponentially from the edge of the disk toward its center, and $\phi^{+}(e^{i\pi/4}cr)$ and $\phi^{-}(e^{-i\pi/4}cr)$ are valid for a circular hole of radius $r_{0}$ in an infinite disk.
The two vector components $a_{1}$, $a_{2}$ of the eigenfunctions $\Psi_{\pm}(kr)$ in Eq. (12) do not depend on the polar angle $\alpha$. Thus, the $S^{2}$ field is uniform in space:
[…]
which describes a fixed semicircle on the unit sphere, beginning with $a_{2}=0$ at the North Pole, hitting at $a_{1}=a_{2}$ the equator and ending with $a_{1}=0$ at the South Pole. For the bulk modes $a_{1,2}$ are real such that $\varphi_{2}-\varphi_{1}=0$ for $E\geq m$ and $\varphi_{2}-\varphi_{1}=\pi$ for $E\leq-m$:
[…]
According to Eq. (14), on the other hand, for the edge modes $a_{1}$ is imaginary and $a_{2}$ is positive. This gives for $k\to e^{\pm i\pi/4}c$ two branches of edge modes with the phases $\varphi_{1}=\pm\pi/2$, $\varphi_{2}=0$ and
[…]
Thus, the bulk and the edge modes are associated with semicircles on the unit sphere, which meet the equator at $(a_{2}/|a_{2}|,0,0)^{T}$ or at $(0,\pm 1,0)^{T}$, respectively, as illustrated by the red curves in Fig. 1. In both cases the EV vanishes due to $\nabla\times{\vec{s}}=0$, while the divergence of $\vec{s}$} with respect to $\bf k$} does not vanishes in the 121-2 plane. Finally, the intensity $(|a_{1}|^{2}+|a_{2}|^{2}|)|\phi^{\pm}|^{2}$ as a function of the radius $r$ decays like $1/r$ for the bulk modes but is strongly localized at the edge for the edge mode.
4.3 Dirac model
Another example is a Dirac Hamiltonian $H_{D}={\vec{h}}_{D}\cdot{\vec{\sigma}}$, which reads with polar coordinates
${\vec{h}}_{D}=(\cos\alpha\ i\partial_{r}-\sin\alpha\frac{1}{r}\ i\partial_{\alpha},\cos\alpha\frac{1}{r}\ i\partial_{\alpha}+\sin\alpha\ i\partial_{r},m).$ (20)
It acts on a circular geometry and has the eigenfunctions (cf. App. A)
[…]
where $C_{n}$ can be expressed as the linear combination of Bessel functions $\phi^{\pm}_{n}$ of Eq. (9). $\rho$, $E$ and $k$ are connected by two conditions (see Eq. (44) in App. A):
$ki\rho=E-m,\ \ k=i\rho(E+m),$ (22)
which implies $\rho^{2}=(m-E)/(m+E)$ and $k^{2}=E^{2}-m^{2}$. The angular quantum numbers are $n=0,\pm 1,\ldots$, and the signs of $k$ and $\rho$ are fixed by the relations (22). Thus, the eigenvalues are degenerate with respect to $n$, since $H_{D}$ is circular invariant. In other words, $n$ is the eigenvalue of the angular momentum operator, and the latter commutes with $H_{D}$.
bulk modes in the upper band:
edge modes:
Figure 3: The 121-2 projection of the $S^{2}$ field $\vec{s}$} of the Dirac model on a circular geometry. Top row: the upper bulk modes for eigenfunctions based on a) $J_{1,2}$, b) $\phi^{+}_{1,2}$, and c) $\phi^{-}_{1,2}$. Bottom row: the edge modes for eigenfunctions based on a) $J_{1,2}$, b) $\phi^{+}_{1,2}$, and c) $\phi^{-}_{1,2}$.
For $E>m$ and E<mE<-m $\rho$ is imaginary and $k=\pm\sqrt{E^{2}-m^{2}}$ is real. On the other hand, for m<E<m-m<E<m $\rho$ is real and $k=\pm\sqrt{E^{2}-m^{2}}$ is imaginary. In particular, for $E=0$ we have either $\rho=1$, k=imk=-im or $\rho=-1$, $k=im$. This again is the analytic continuation $k\to ic$ for bulk to edge modes with $0\leq c^{2}\leq m^{2}$, illustrated in Fig. 2b. Thus, from Eq. (9) we obtain edge modes with eigenvalues $E_{c}=\pm\sqrt{m^{2}-c^{2}}$, which either grow or decay exponentially with $r>0$ on the scale $1/c=1/\sqrt{m^{2}-E^{2}}$ for m<E<m-m<E<m. This scale diverges as $E$ approach the spectral boundaries $\pm m$ of the edge modes. On the other hand, the wavenumber of the bulk states vanishes as $k=\sqrt{E^{2}-m^{2}}$ for $E^{2}\sim m^{2}$. Therefore, $E=\pm m$ are singular points in the spectrum, where the edge modes become uniformly extended over the entire 2D system. This is reminiscent of a localization-delocalization (or Anderson) transition in disordered systems and reflects the critical energy dependence of the edge modes and its transition to a bulk mode. These results suggest the introduction of an index $b$ in the wavefunction of Eq. (21)
[…]
where $b$ is either the band index $b=\uparrow,\downarrow$ for the upper and the lower band, or $b=\pm$ for the energies $E=\pm\sqrt{m^{2}-c^{2}}$ of the edge modes. This gives for the $S^{2}$ field
[…]
The 121-2 projection is visualized for several eigenfunctions in Fig. 3. The phase shift $\eta_{b}$ is the angle between the radial vector $\bf r$} and the projected vector $(s_{1},s_{2})^{T}$. Comparing this expression with Eq. (4), we get the phase relation between the eigenfunctions and the $S^{2}$ field as
$\varphi_{2}-\varphi_{1}=\alpha+\eta_{b}.$ (25)
Besides the different energies for the bands of the bulk and for the edge modes, also the parameter $\rho_{b}(m,k)$ distinguishes the different modes. With the band energies $E=\pm\sqrt{m^{2}+k^{2}}$ we get from the relations in Eq. (22) for the bulk an imaginary parameter
$\rho_{\uparrow}=-i\sqrt{\frac{\sqrt{m^{2}+k^{2}}-m}{\sqrt{m^{2}+k^{2}}+m}}=1/\rho_{\downarrow}.$ (26)
On the other hand, for the edge modes we have a real parameter
$\rho_{+}=\sqrt{\frac{m-\sqrt{m^{2}-c^{2}}}{m+\sqrt{m^{2}-c^{2}}}}=1/\rho_{-}.$ (27)
This entails that under $m\to-m$ we have $\rho_{\uparrow}\leftrightarrow-\rho_{\downarrow}$ and $\rho_{+}\leftrightarrow\rho_{-}$. The behavior of the $\rho_{b}(k)$ is presented in Fig. 5. Since ${\rm arg}(\rho_{b})$ is independent of the values of $m$ as well as $k$ and $c$, respectively, it can be understood as a topological number, analogous to the Chern number, which is associated with the band:
${\rm arg}(\rho_{b})=\cases{-\pi/2&$b=\uparrow$\cr\pi/2&$b=\downarrow$\cr 0&$b=\pm$\cr}.$ (28)
The phase shift $\eta_{b}$ in ${\vec{s}}_{b}$ of Eq. (24) has different values for the two bands and for the edge modes due to $\eta_{b}={\rm arg}(C_{n}^{*}C_{n+1})+{\rm arg}(\rho_{b})$. $\sin\eta_{b}(r)$ is plotted in Fig. 4 for bulk and edge modes and for different eigenfunctions.
Figure 4: $\sin\eta_{b}$ as a function of the radius $r$ for the Dirac model with $\phi^{+}_{1,2}$ (left) and $\phi^{-}_{1,2}$ (right).
There are transitions from bulk to edge modes at $E=m$ and from edge modes to bulk modes at E=mE=-m. They are associated with a change of ${\rm arg}(\rho_{b})$ or by a sign change of $\sin\eta_{b}$. The latter appears in the EV as
${\cal V}=\int_{0}^{2\pi}{\vec{s}}\cdot\frac{d{\bf r}}{d\alpha}d\alpha=4\pi r_{0}\frac{|C_{n}^{*}C_{n+1}\rho_{b}|}{|C_{n}|^{2}+|\rho_{b}C_{n+1}|^{2}}\sin\eta_{b}(r_{0})$ (29)
after an integration with respect to the edge of a circular hole with radius $r_{0}$. EV is positive in the upper band and negative in the lower band (cf. Fig. 4). Since $\eta_{b}$ depends on the radius, the EV can change with the position of the edge and the boundary conditions.
Figure 5: The parameter $\rho_{b}(k)$ and the eigenvalues $E$ for the bulk (a) and for the edge modes (b) of the Dirac model with $m=1$. $k=0$ and $c=0$ are the critical points $E=\pm 1$ for the localization-delocalization transition as well as as a significant change of the parameter $\rho_{b}$ (cf. Eqs. (26), (27)).
Figure 6: Radial intensity distribution $I(r)$ for the Dirac model on a disk with radius $r_{0}=30$ for bulk (blue) and edge modes (yellow). The eigenfunctions are based on a) $J_{1,2}$, b) $\phi^{+}_{1,2}$, and c) $\phi^{-}_{1,2}$.
$\rho_{\downarrow}$ and $\rho_{-}$ diverge for $k\to 0$, while $\rho_{\uparrow}$ and $\rho_{+}$ vanish in this limit (cf. Fig. 5), reflecting the qualitative change of the wavefunctions from edge-localized functions to circular waves at $E=\pm m$. At the critical point E=mE=-m the $S^{2}$ vector reaches the South Pole. On the other hand, at the critical point $E=m$ the $S^{2}$ vector arrives at the North Pole. The trajectories of ${\vec{s}}(\alpha)$ at fixed $r$ are indicated by the blue circles in Fig. 1. Finally, the intensity $I=|C_{n}|^{2}+|\rho_{b}C_{n+1}|^{2}$ is plotted in Fig. 6 for different eigenfunctions. In all three cases the main difference is the significant weight of the edge mode intensity either at the center of the disk or at its boundary.
5 Discussion and conclusions
After having solved the eigenvalue equation for the bulk modes, the analytic continuation of the real wavenumber $k$ into the complex plane provides the edge modes. This approach yields a spectral separation of bulk and edge modes. On the other hand, there is no complete spatial separation of these two types of modes, since the edge modes extend into the bulk with an exponential decay on the scale $\xi$:
$\xi=\cases{1/\sqrt{|E|}&single Laplacian\cr\sqrt{2}/|(m^{2}-E^{2})^{1/4}|&two-band Laplacian model\cr 1/\sqrt{m^{2}-E^{2}}&Dirac model\cr}.$ (30)
This means that the edge modes are two-dimensional, which cannot be described as solutions of a one-dimensional equation. However, they do not spread over the entire system but decay exponentially from the edge. In contrast, the bulk modes decay like $1/\sqrt{r}$.
By varying $E$ we can scan continuously through the bulk and the edge spectrum. There are bulk-edge transitions at $E=\pm m$, representing transitions between localized edge modes and delocalized bulk modes. They are accompanied by a qualitative change for the Dirac model, since the index ${\rm arg}(\rho_{b})$ and the sign of $\sin\eta_{b}$ change according to Eq. (28) and Fig. 4. For the two-band Laplacian model the phase difference $\Delta\varphi:=\varphi_{2}-\varphi_{1}$ switches from $\Delta\varphi=0$ for the upper band and $\Delta\varphi=\pi$ for the lower band to $\Delta\varphi=\pm\pi/2$ for the two branches of the edge modes. This indicates that the bulk and the edge modes are not only characterized by their spatial decay but also by the phase difference $\Delta\varphi$ in the $S^{2}$ field $\vec{s}$}.
5.1 Boundary conditions
Next, we introduce boundary conditions for the eigenfunctions of the Hamiltonian, whose purpose is two-fold, namely (i) to get unique solutions for the physical system (i.e., there can be no ambiguity in terms of physical results) and (ii) to obtain self-adjoint operators for a given Hilbert space: Although we have found the solution of $H\Psi_{E}=E\Psi_{E}$, to diagonalize the Hamiltonian matrix we still need that the inner product $(\Psi_{E},\Psi_{E^{\prime}})=0$ for $E\neq E^{\prime}$. This is equivalent with the request that $H$ is self-adjoint. For instance, the Laplacian $\Delta$ is diagonal if the boundary terms in
$(\Psi_{E},\Delta\Psi_{E^{\prime}})=(\Delta\Psi_{E},\Psi_{E^{\prime}})+{\rm boundary\ terms}$ (31)
vanish due to appropriate boundary conditions. We can also directly obtain such a relation by using the fact that we operate in the eigenbasis of the Laplacian for a disk geometry with radius $R$, using Bessel functions. Thus, the Hilbert space is spanned by Bessel functions on the interval $0\leq r\leq R$, where the inner product reads
$(\psi_{kn},\psi_{qn^{\prime}})=\int_{0}^{2\pi}e^{i(n^{\prime}-n)\alpha}d\alpha\int_{0}^{R}\phi_{n}(kr)\phi_{n^{\prime}}(qr)rdr=\delta_{nn^{\prime}}\int_{0}^{R}\phi_{n}(kr)\phi_{n}(qr)rdr$ (32)
with $\phi_{n}=J_{n}+AY_{n}$ and a real coefficient $A$. For a self-adjoint Hamiltonian this relation requires the orthogonality of the eigenfunctions; i.e., $\delta_{kq}$ for the integral on the right-hand side. The latter can be obtained with proper boundary conditions (rule 11.4.3 in Ref. [28]) since
$\int_{0}^{R}\phi_{n}(kr)\phi_{n}(qr)rdr=\frac{R}{q^{2}-k^{2}}\left[k\phi_{n}(qR)\phi^{\prime}_{n}(kR)-q\phi_{n}(kR)\phi^{\prime}_{n}(qR)\right].$ (33)
Then the Bessel functions are orthogonal, for instance, for the zeros of $\phi_{n}$ or of $\phi^{\prime}_{n}$. In other words, there is a sequence $\{k_{n,l}\}_{l}$ of wave numbers with $\phi_{n}(k_{n,l}R)=0$ that provides the orthogonal eigenfunctions of the Laplacian. Another sequence $\{k^{\prime}_{n,l}\}$ is obtained from $\phi^{\prime}_{n}(k^{\prime}_{n,l}R)=0$ as boundary condition. Corresponding boundary conditions for the Dirac operator are obtained for $\{k_{nl}\}$, which are solutions of
$A=-\frac{J_{n}(k_{nl}R)}{Y_{n}(k_{nl}R)}=-\frac{J_{n+1}(k_{nl}R)}{Y_{n+1}(k_{nl}R)}$ (34)
and imply $\phi_{n}(k_{nl}R)=\phi_{n+1}(k_{nl}R)=0$.
In the next step we apply boundary conditions to the edge modes on a disk with radius $R$. Starting from general bulk eigenfunctions of the Laplacian, we perform the analytic continuation of $k\to\pm ic$. Assuming bounded solutions, we are enforced to use $I_{n}(cr)$ because $K_{n}(cr)$ is singular at $r=0$. Since $I_{n}(cr)$ is a monotonic increasing function, only mixed boundary conditions can be satisfied: $I_{n}(cR)-bcI^{\prime}_{n}(cR)=0$ with a real parameter $0<b<\infty$, which can also be written with $z=cR$ as
$\frac{d\log I_{n}(z)}{dz}=\frac{I^{\prime}_{n}(z)}{I_{n}(z)}=\frac{1}{bc}=\frac{R}{bz}.$ (35)
Since $d\log(I_{n}(z))/dz$ increases monotonically, while the right-hand side decreases monotonically, there is only a single solution $c_{n}$ for this boundary condition at fixed $n$. Thus, after fixing the boundary conditions we have a point spectrum $\pm\sqrt{m^{2}-c_{n}^{2}}$ on the circle in Fig. 2b, and there is no continuous transition from the edge to the bulk spectrum.
5.2 Properties of the $S^{2}$ field
The $S^{2}$ real vector field $\vec{s}$} characterizes the gauge-invariant properties of the eigenfunctions. We briefly summarize its behavior, visualized by the unit sphere in Fig. 1. Under the variation of the polar coordinates $(r,\alpha)$ the trajectory of the unit vector depends strongly on the model. For the two-band Laplace model there is no $\alpha$ dependence, such that the trajectory follows the meridians. The meridian trajectory jumps by an angle of $\pi/2$ when we switch from bulk to edge modes, as stated in Eqs. (18) and (19). This reflects the fact that the model is topological trivial. In contrast, the trajectories of Dirac Hamiltonian in Eq. (24) cover the entire sphere, where a variation of $r$ follows also the meridians, while a variation of $\alpha$ for $0\leq\alpha<2\pi$ gives a full circle at fixed latitude, reflecting that the winding number of $\vec{s}$} is 1. Going from bulk to edge states at fixed $\alpha$ results in a jump of the meridian, similar to the two-band Laplace model. Thus, the winding number of $\vec{s}$} can be used for the characterization of the Hamiltonian. This should also be applicable to Hamiltonians with more than two bands.
5.3 Comparison with the TBEC approach
The TBEC for a straight (infinite) edge was studied for the 2D Dirac model [20] and for the extended hydrodynamic model [22], using the relation between the Chern number $C_{+}$ of the bulk modes in the upper band. The TBEC holds if the number $n_{e}$ of edge modes is equal to $C_{+}$. For the translation-invariant 2D Dirac operator the Chern number reads
$C_{+}=\frac{1}{2\pi i}\int_{\mathbb{R}^{2}}{\rm tr}(P_{+}[\partial_{k_{x}}P_{+},\partial_{k_{y}}P_{+}])dk_{x}dk_{y}=1,$ (36)
where $P_{+}$ is the eigenprojection for the upper band. Depending on the boundary conditions, though, the TBEC does not always hold, since $n_{e}=2,3$ for some boundary conditions [20]. The origin of this violation of the TBEC has been associated with the unbounded spectrum of the translation-invariant 2D Dirac operator. In contrast, the edge is finite for the circular-invariant 2D Dirac operator, implying that also its spectrum is bounded. The difference between the straight edge and the circular edge can be seen already in the spectrum before employing boundary conditions: for an edge along the $x$-axis we have $E(k_{x},c)=\pm\sqrt{m^{2}+k_{x}^{2}-c^{2}}$ with $\infty<k_{x}<\infty$, $c^{2}\leq m^{2}+k_{x}^{2}$, while $E(n,c)=\pm\sqrt{m^{2}-c^{2}}$ with $c^{2}\leq m^{2}$ for the circular edge is bounded. Another example for the violation of the TBEC appears for the two-band Laplacian, where $C_{+}=0$. As discussed in Sect. 4.2 and in Sect. 5.1, there are edge modes. It should be noted that the ABEC does not lead to any conflict in this case. However, the counting of the number of edge modes that merge with the bulk modes through the Chern number is not possible here.
5.4 Robustness
An important question concerns the robustness of the edge modes when we replace the uniform mass $m$ by a spatially varying mass $m(r,\alpha)$. In the circular-symmetric case $m(r)$, a sign change of the mass creates an additional edge mode with a skyrmion-like wavefunction [7]. Moreover, with a positive $m(r,\alpha)={\bar{m}}+\delta m(r,\alpha)$ we can break the circular symmetry. If $\delta m(r,\alpha)/{\bar{m}}\ll 1$, the robustness of the edge modes can be analyzed within perturbation theory. The (degenerate) perturbation expansion in powers of $\delta m(r,\alpha)$ would provide a stability analysis of the edge modes. A thorough study of this perturbation approach exceeds the scope of this paper and should be left for a separate project in the future.
5.5 Conclusions
As a summary of the results of the two-band examples, we found the bulk eigenfunctions of circular symmetric two-band models. They are of the form
[…]
with $k$ real and $\phi_{1,2}$ complex. Physical properties are obtained from the $S^{2}$ field $\vec{s}$}
[…]
The 121-2 projection of the $S^{2}$ field identifies the phase difference of the two components of the eigenfunctions. An analytic continuation of $k$ yields the corresponding expressions for the edge modes. From these results we conclude that the ABEC offers a systematic approach through an analytic continuation of the wavenumber for the description of two-band Hamiltonians with edges. Although we have focused here on a circular geometry for simplicity, the concept can be extended to other geometries with finite edges.
Acknowledgment
I am grateful to the anonymous Referee for bringing the works of Refs. [8, 20, 22] to my attention.
Appendix A Eigenfunctions of the Dirac Hamiltonian
With the ansatz
[…]
we can write for the eigenvalue problem $H_{D}\Phi_{E,n}=E\Phi_{E,n}$ the equation
[…]
which enables us to eliminate $e^{in\alpha}$ and $e^{i(n+1)\alpha}$, respectively, on both sides of the second equation:
[…]
Now we consider that $C_{n}$ is either the Bessel function $J_{n}$, $Y_{n}$ or a linear combination of these two functions and write $f_{n}(kr)=C_{n}(kr)$ and $g_{n}(kr)=\rho C_{n+1}(kr)$. Then we employ the recurrence relations of the Bessel functions [28]
[…]
and
[…]
With these relations we obtain from Eq. (39)
[…]
or the eigenvalue equation
[…]
which gives the relations
$k\rho=i(m-E),\ \ k=i(E+m)\rho.$ (44)
This determines the parameter $\rho$ in $g_{n}$. Thus, we have $k^{2}=E^{2}-m^{2}$ and $\rho^{2}=(m-E)/(m+E)$, both are independent of $n$. The eigenfunction in Eq. (37) becomes
[…]
Appendix B Analytic continuation of Bessel functions
The analytic continuation $k\to ic$ with a real $c$ yields for the Bessel functions [28]
$J_{n}(kr)\to J_{n}(icr)=e^{i\pi n/2}I_{n}(cr)\ ,\ \ H_{n}^{(1)}(kr)\to H_{n}^{(1)}(icr)=-\frac{2i}{\pi}e^{-i\pi n/2}K_{n}(cr)$ (46)
with $H_{n}^{(1)}(kr)=J_{n}(kr)+iY_{n}(kr)$, which implies
$J_{n}(icr)^{*}J_{n+1}(icr)=iI_{n}(cr)I_{n+1}(cr)$ (47)
and
$H_{n}^{(1)}(icr)^{*}H_{n+1}^{(1)}(icr)=-\frac{4i}{\pi^{2}}K_{n}(cr)K_{n+1}(cr).$ (48)